# DESY-25-105

### v1

@@ Abstract
We propose an algebro-geometric interpretation of the Schur and Macdonald indices of four-dimensional $\mathcal{N}=2$ superconformal field theories (SCFTs). We conjecture that there exists an affine scheme $X$ such that the Hilbert series of the (appropriately-graded) arc space of its polynomial ring $J_{\infty}(\mathbb{C}[X])$ encodes the indices. Distinct local descriptions of a (singular) point correspond to distinct choices of $X$, giving rise to a family of $\mathcal{N}=2$ SCFTs each without a Higgs branch. These local descriptions directly translate into nilpotency relations in the operator product expansions. We test our conjecture across a variety of (generalized) Argyres–Douglas (AD) theories.
@@ 1 Introduction
One of the key insights gained over the past few decades in quantum field theory (QFT) is that the conventional Lagrangian and path-integral formulation does not capture the full landscape of consistent quantum field theories. This realization stems from the existence of a wide variety of inherently strongly coupled theories, often referred to as “non-Lagrangian” QFTs, which defy a standard Lagrangian description. In the conventional approach, a QFT is defined by specifying a set of fields and their interactions encoded in a Lagrangian. However, such a framework proves inadequate for describing these exotic non-Lagrangian theories. This raises a natural and pressing question: is there a more general formulation of QFT that can encompass both conventional and non-Lagrangian theories on equal footing?
@@
[…]
Let $\mathcal{T}$ denote some 4d $\mathcal{N}=2$ SCFT. We propose that there exists an affine algebraic scheme $X$ associated with $\mathcal{T}$ such that the underlying variety of $X$ is the Higgs branch of $\mathcal{T}$:
@@
As the name suggests, the arc space $J_{\infty}(Y)$ encodes all local paths (arcs) passing through the space $Y$. This allows us to further study the local or singular properties of $Y$. We review the definitions and properties of the arc space (and the, related, jet schemes) in Section 2.1. The arc space has a natural bi-grading, and one can determine the Hilbert series of the (polynomial ring over the) arc space with respect to this bi-grading. One obtains a polynomial:
@@
Here, the fugacity $p$ denotes the grading of the monomials that exists before going to the arc space, and $q$ denotes the arc space grading. One can further refine the Hilbert series by additional, “flavor”, symmetries, if they are present in $\mathcal{T}$.
@@
Furthermore, we can construct an associated graded algebra $\mathrm{gr}(J_{\infty}(R))$ to $J_{\infty}(R)=J_{\infty}(\mathbb{C}[X])$, and the Hilbert series of that graded algebra is a multi-graded polynomial, for which a certain specialization provides the Macdonald index [49] of $\mathcal{T}$:
$\boxed{I_{\textrm{Mac}}(q,T)=\operatorname{HS}_{p_{i},q}\mathrm{gr}(J_{\infty}(R))\Big{|}_{p_{i}\to q^{d_{i}}T^{w_{i}}}}\,.$ (1.5)
We will explain this expression in detail in Section 4. This is essentially a geometric analogue of the procedure in [74]. The Macdonald index counts the same set of states as the Schur index, but with refined information of $SU(2)_{R}$ grading.
While it may be tempting to claim that all affine schemes give rise to (the indices of) a 4d theory, we will see later that not all affine schemes have arc spaces that can be associated to indices of consistent 4d $\mathcal{N}=2$ SCFTs. Instead, we might conjecture that there exists a restricted class of affine schemes, specified by currently unknown algebro-geometric conditions, such that for each such affine scheme there is a corresponding 4d $\mathcal{N}=2$ superconformal field theory $\mathcal{T}_{X}$ modulo exactly marginal deformations (and also possible global structures): [^2] Here the choice of global structures, such as the 1-form symmetry or discrete theta terms may not specified. Therefore, at best it maps to a relative theory.
[…]
@@
$\begin{array}[]{c|c}\mathcal{T}_{X}&X\\ \hline\cr(A_{1},A_{2n})&x^{n+1}=0\\ (A_{2},A_{3})&x^{2}-y^{3}=0\ ,\quad xy=0\\ (A_{2},A_{4})&x^{2}-y^{3}=0\ ,\quad xy^{2}=0\\ (A_{2},A_{3m})&x^{m-1}(x^{2}-y^{3})=0\ ,\quad x^{m-2}y(x^{2}-y^{3})=0\quad(m\geq 2)\\ (A_{2},A_{3m+1})&x^{m-1}(x^{2}-y^{3})=0\ ,\quad x^{m-2}y^{2}(x^{2}-y^{3})=0\quad(m\geq 2)\\ (A_{3},A_{4})&xz+y^{2}+z^{3}=0,\quad xy+yz^{2}=0,\quad x^{2}+y^{2}z+z^{4}=0\\ (A_{1},D_{2n+1})&w=x^{2}+y^{2}+z^{2}\ ,\quad xw^{n}=yw^{n}=zw^{n}=0\\ (A_{1},A_{2n-1})&xy+z^{n}-tz^{n-2}=0\,\quad xt=0\,,\quad yt=0\,,\quad z(xy+z^{n})=0\end{array}$ (1.16)
@@
and the algebraic variety underlying the associated scheme, called the associated variety, can be obtained by reducing the scheme, dropping the nilpotents. [^4] The spectrum of a ring $\operatorname{Spec}(R)$ is the set of all prime ideals of $R$, whereas $\operatorname{Specm}(R)$ denotes the set of all maximal ideals of $R$. Recall that not all prime ideals are maximal. This is conjectured to be isomorphic to the Higgs branch of the 4d SCFT [17]:
$(X_{V})_{\text{red}}=\operatorname{Specm}(R_{V})=\mathcal{M}_{H}(\mathcal{T})\,.$ (1.21)
Since the spectrum of $R_{V}$ encodes decoupling relations in the OPEs of certain protected operators, and is conjectured to reduce as a variety to the Higgs branch, we are motivated to ask if there exists a ring $R$ with $X=\operatorname{Spec}(R)$, which might or might not be $X_{V}$, that captures all decoupling relations necessary to reconstruct the index.
@@
Such examples are drawn from the $(A_{k-1},A_{N-1})$ Argyres–Douglas SCFTs with $(k,N)=1$. For example, Zhu’s $C_{2}$-algebra for the $(A_{1},A_{2n})$ Argyres–Douglas theory is given as [17]
@@
Our goal is to ask the opposite direction. For a given affine scheme $X$, when do we have a 4d $\mathcal{N}=2$ superconformal theory whose Hilbert series of $J_{\infty}(\mathbb{C}[X])$ reproduces the index? We start with various descriptions of singularity, identical as an affine variety but different as an affine scheme. We find that whenever we obtain a valid index of a 4d SCFT, in most cases, our affine scheme $X$ is identical to the associated scheme $X_{V}$, which is identical to the non-reduced spectrum of the Zhu’s $C_{2}$-algebra $R_{V}$ of the associated VOA $V$ [15, 17]. However, there is a wrinkle in the case for the theories with the Higgs branch. When $V$ is given by a simple affine Lie algebra, we find that $X$ is not identical to the associated scheme $X_{V}$. So the arc space of Zhu’s $C_{2}$-algebra $R_{V}$ does not reproduce a sensible Schur index. To fix this, we need to include an additional coordinate, which has an interpretation as the stress-energy tensor. We discuss these examples in Section 5.
@@ 2.1 Arc spaces: a primer
The jet scheme $J_{n}(X)$ admits a natural bi-grading defined by
@@
Practically, this involves enumerating the infinitely many coordinates and polynomials in the ideal for the arc space of a given scheme $X$. Intuitively, the arc space captures formal smooth paths (or arcs) based at points of $X$, with the bi-degree keeping track of derivative order. For example, the first jet scheme $J_{1}(X)$ can be identified with the total space of the tangent bundle of $X$.
@@ 2.2 Vanishing short multiplets in Argyres–Douglas theories
This phenomenon can be understood from the perspective of the associated VOA. The VOA corresponding to the $(A_{1},A_{2n})$ theory is the Virasoro minimal model $M(2,2n+3)$, which contains a null state of the form $((L_{-2})^{n}+\cdots)\Omega$ at level $2n$, where $\Omega$ denotes the vacuum state. This reflects the vanishing of the $\widehat{\mathcal{C}}_{n(\frac{n}{2},\frac{n}{2})}$ multiplet in the physical theory.
@@
where $\operatorname{PE}$ denotes the standard plethystic exponential. The term $\frac{q^{2}}{1-q}$ arises from the contribution of the stress tensor, while the term $-\frac{q^{2n+2}}{1-q}$ reflects the absence of the $\widehat{\mathcal{C}}_{n(\frac{n}{2},\frac{n}{2})}$ multiplet. [^9] The Schur index alone does not uniquely determine which short multiplet is absent; this requires additional input from OPE selection rules [66, 70, 62, 56]. In some cases, further information can be obtained by studying the Macdonald or full superconformal index, combined with character decomposition of short multiplets [2, 75].
@@
This motivates our affine scheme $X$ to be given as follows:
$(A_{2},A_{3}):$ $\quad x^{2}-y^{3}=0\,,\ xy=0\,,$ (2.18a)
$(A_{2},A_{4}):$ $\quad x^{2}-y^{3}=0\,,\ xy^{2}=0\,,$ (2.18b)
$(A_{2},A_{3m}):$ $\quad x^{m-1}(x^{2}-y^{3})=0\ ,\ x^{m-2}y(x^{2}-y^{3})=0\,,\quad\,(m\geq 2)$ (2.18c)
$(A_{2},A_{3m+1}):$ $\quad x^{m-1}(x^{2}-y^{3})=0\ ,\ x^{m-2}y^{2}(x^{2}-y^{3})=0\,.\quad(m\geq 2)$ (2.18d)
Here $x$ and $y$ comes from $W$ and $T$ respectively. In each case the scheme reduces to a point when reduced to an affine variety. But, as a scheme, it has a more intricate singularity structure, which captures the physics of the indices.
We can further generalize this to an arbitrary $(A_{k-1},A_{N-1})$ theory. The associated VOA is a $W_{k}$-algebra, containing $k-1$ generators of spins $2,3,\cdots,k$. The Schur index (or the vacuum character of the $W_{k}$-algebra) is given by
@@
$\begin{split}q^{6}:&\quad W_{4}W_{2}+(W_{3})^{2}+(W_{2})^{3}=xz+y^{2}+z^{3}=0\,,\\ q^{7}:&\quad W_{4}W_{3}+W_{3}(W_{2})^{2}=xy+yz^{2}=0\,,\\ q^{8}:&\quad(W_{4})^{2}+(W_{3})^{2}W_{2}+(W_{2})^{4}+(W_{2})^{2}W_{4}=x^{2}+y^{2}z+z^{4}+z^{2}x=0\,,\end{split}$ (2.20)
where we ignored the coefficients. Writing $(W_{2},W_{3},W_{4})\to(z,y,x)$ will give the desired relations for the point-like singularity in three dimensions. It is possible to get the same relations from $F=z^{5}+z^{3}x+zx^{2}+z^{2}y^{2}+y^{2}x$ and taking derivatives with respect to $x,y,z$.
More generally, it is conjectured that for the Argyres–Douglas theories of type $(A_{k-1},A_{n-1})$ with $(k,n)=1$, Zhu’s $C_{2}$-algebra $R_{V}$ is given by a simple Jacobi algebra of the form
$R_{V}=\mathbb{C}[w_{1},\ldots,w_{\ell}]\Big{/}\left(\frac{\partial F}{\partial w_{1}},\ldots\frac{\partial F}{\partial w_{\ell}}\right)\,,$ (2.21)
where $F=F(w_{1},\ldots w_{\ell})$ is the sum of monomials of fixed degree and the degrees for $w_{i}$ are given as $i+1$ [80]. If we use their conjecture and choose $\mathbb{C}[X]=R_{V}$ as above to compute the Hilbert series of its arc space, we find the correct Schur index for the $(A_{k-1},A_{n-1})$ theory. This supports the proposal of [80] and identification of $X$ with the associated scheme $X_{V}=\operatorname{Spec}R_{V}$ for the $(A_{k-1},A_{n-1})$ theory.

### v2

@@ Abstract
We propose an algebro-geometric interpretation of the Schur and Macdonald indices of four-dimensional $\mathcal{N}=2$ superconformal field theories (SCFTs). We conjecture that there exists an affine scheme $X$, which reduces to the Higgs branch as a variety, such that the Hilbert series of the (appropriately-graded) arc space of its polynomial ring $J_{\infty}(\mathbb{C}[X])$ encodes the indices. Distinct local descriptions of a (singular) point correspond to distinct choices of $X$, giving rise to families of $\mathcal{N}=2$ SCFTs each without a Higgs branch. These local descriptions directly translate into nilpotency relations in the operator product expansions. We test our conjecture across a variety of (generalized) Argyres–Douglas theories.
@@ 1 Introduction
One of the key insights gained over the past few decades in quantum field theory (QFT) is that the conventional Lagrangian and path-integral formulation does not capture the full landscape of consistent quantum field theories. This realization stems from the existence of a wide variety of inherently strongly-coupled theories, often referred to as “non-Lagrangian” QFTs, which defy a standard Lagrangian description. In the conventional approach, a QFT is defined by specifying a set of fields and their interactions encoded in a Lagrangian. However, such a framework proves inadequate for describing these exotic non-Lagrangian theories. This raises a natural and pressing question: is there a more general formulation of QFT that can encompass both conventional and non-Lagrangian theories on equal footing?
@@
Motivated by these observations, we explore whether the decoupling phenomenon can be elevated to a defining characteristic of a specific CFT. In this paper, we conjecture that for 4d $\mathcal{N}=2$ SCFTs, particular limits of the supersymmetric partition function (namely, the Schur or Macdonald index) can be entirely constructed by a decoupling or vanishing condition in the spectrum. We propose a translation of this decoupling condition into algebraic geometry in terms of nilpotency relations, such as $x^{n}=0$. [^1] This idea is motivated by the behavior of operator product expansions (OPEs), where certain components vanish when the same primary operator is multiplied several times. Such a phenomenon is realized, for example, in the $(A_{1},A_{2n})$ Argyres–Douglas theory [2], where the short multiplet $\widehat{\mathcal{C}}_{n(\frac{n}{2},\frac{n}{2})}$, which generically appears in the OPE of the form $(\widehat{\mathcal{C}}_{0(0,0)})^{n+1}$, is absent. While the equation $x^{n}=0$ naively describes the point $x=0$, viewing it in the context of polynomial rings reveals the richer structure imposed by the nilpotent relation. This distinction reflects the essential difference between a scheme and a variety in algebraic geometry, and it plays a central role in our proposal.
Let $\mathcal{T}$ denote some 4d $\mathcal{N}=2$ SCFT. We propose that there exists an affine algebraic scheme $X$, necessarily bifiltered to capture the structure of the superconformal algebra, associated with $\mathcal{T}$ such that the underlying variety of $X$ is the Higgs branch of $\mathcal{T}$:
@@
As the name suggests, the arc space $J_{\infty}(Y)$ encodes all local paths (arcs) passing through the space $Y$. This allows us to further study the local or singular properties of $Y$. We review the definitions and properties of the arc space (and the, related, jet schemes) in Section 2.1. Given a grading on $Y$, which we refer to via the $q$-degree, the arc space has a natural bigrading, and one can determine the Hilbert series of the (polynomial ring over the) arc space with respect to this bigrading. One obtains a polynomial:
@@
Here, the fugacity $q$ denotes the grading of the monomials that exists before going to the arc space, and $p$ denotes the arc space grading. One can further refine the Hilbert series by additional, “flavor”, symmetries, if they are present in $\mathcal{T}$.
@@
More generally, the $q$-grading of $X$ extends to part of a bifiltration structure on $X$, where the second part of the bifiltration captures the $\mathfrak{su}(2)_{R}$ symmetry of the $\mathcal{N}=2$ superconformal algebra. Keeping track of the full bifiltration, we can then construct an associated graded ring $\operatorname{gr}(J_{\infty}(R))$ to $J_{\infty}(R)=J_{\infty}(\mathbb{C}[X])$, which is trigraded. The Hilbert series is then a polynomial in three fugacities, $q$, $p$, and $T$, for which a certain specialization provides the Macdonald index [49] of $\mathcal{T}$:
$\boxed{I_{\textrm{Mac}}(q,T)=\operatorname{HS}_{p,q,T}\left(\mathrm{gr}(J_{\infty}(R))\right)\Big|_{p\to q}}\,.$ (1.5)
The $T\rightarrow 1$ limit, which recovers the Schur index, corresponds to throwing up the information about the full bifiltration and keeping only the $q$-grading. We will explain this expression in detail in Section 4, and we also refer to [60]. This is essentially a geometric analogue of the procedure in [77]. The Macdonald index counts the same set of states as the Schur index, but with refined information of $\mathfrak{su}(2)_{R}$ grading.
While it may be tempting to claim that all affine schemes give rise to (the indices of) a 4d theory, we will see later that not all affine schemes have arc spaces that can be associated to indices of consistent 4d $\mathcal{N}=2$ SCFTs. Instead, we might conjecture that there exists a restricted class of affine schemes, specified by currently unknown algebro-geometric conditions, [^2] In [60], the algebro-geometric conditions are worked out for the $(A_{k-1},A_{N-1})$ Argyres–Douglas theories with $\gcd(k,N)=1$. such that for each such affine scheme there is a corresponding 4d $\mathcal{N}=2$ superconformal field theory $\mathcal{T}_{X}$ modulo exactly marginal deformations (and also possible global structures): [^3] Here the choice of global structures, such as the 1-form symmetry or discrete theta terms may not be specified. Therefore, at best it maps to a relative theory.
$\{\textrm{Affine schemes}^{\prime}\}$ $\to\{\textrm{4d }\mathcal{N}=2~\textrm{(relative) SCFTs (up to marginal deformations)}\}\,.$ (1.6)
@@
$\begin{array}[]{c|c}\mathcal{T}_{X}&X\\ \hline\cr(A_{1},A_{2n})&x^{n+1}=0\\ (A_{2},A_{3})&x^{2}+y^{3}=0\ ,\quad xy=0\\ (A_{2},A_{4})&x^{2}+y^{3}=0\ ,\quad xy^{2}=0\\ (A_{2},A_{6})&x^{3}+4xy^{3}=0\ ,\quad 3x^{2}y+y^{4}=0\\ (A_{2},A_{7})&y^{5}+5y^{2}x^{2}=0\ ,\quad 2x^{3}+3xy^{3}=0\\ (A_{3},A_{4})&4xz+y^{2}+8z^{3}=0,\quad xy+4yz^{2}=0,\quad x^{2}+4y^{2}z+12xz^{2}+22z^{4}=0\\ (A_{1},D_{2n+1})&w=x^{2}+y^{2}+z^{2}\ ,\quad xw^{n}=yw^{n}=zw^{n}=0\\ (A_{1},A_{5})&xy+z^{3}-tz=0\,\quad xt=0\,,\quad yt=0\,,\quad z(xy+z^{3})=0\end{array}$ (1.16)
@@
and the algebraic variety underlying the associated scheme, called the associated variety, can be obtained by reducing the scheme, dropping the nilpotents. This is conjectured to be isomorphic to the Higgs branch of the 4d SCFT [17]:
$(X_{V})_{\text{red}}=\mathcal{M}_{H}(\mathcal{T})\,.$ (1.21)
Since the spectrum of $R_{V}$ encodes decoupling relations in the OPEs of certain protected operators, and is conjectured to reduce as a variety to the Higgs branch, we are motivated to ask if there exists a ring $R$ with $X=\operatorname{Spec}(R)$, which might or might not be $X_{V}$, that captures all decoupling relations necessary to reconstruct the index. [^5] We note that if the VOA is classically-free [80], then the conjecture in equation (1.4) holds for $X=X_{V}$. It has been conjectured (see [72, 73]) that all VOAs that are associated to 4d $\mathcal{N}=2$ SCFTs are classically-free. We thank Chris Beem and Leonardo Rastelli for discussions on this point.
@@
Such examples are drawn from the $(A_{k-1},A_{N-1})$ Argyres–Douglas SCFTs with $\gcd(k,N)=1$. For example, Zhu’s $C_{2}$-algebra for the $(A_{1},A_{2n})$ Argyres–Douglas theory is given as [17]
@@
Our goal is to ask the opposite direction. For a given affine scheme $X$, when do we have a 4d $\mathcal{N}=2$ superconformal theory whose Hilbert series of $J_{\infty}(\mathbb{C}[X])$ reproduces the index? We start with various descriptions of singularity, identical as an affine variety but different as an affine scheme. We find that whenever we obtain a valid index of a 4d SCFT, in most cases, our affine scheme $X$ is identical to the associated scheme $X_{V}$, which is identical to the non-reduced spectrum of the Zhu’s $C_{2}$-algebra $R_{V}$ of the associated VOA $V$ [15, 17]. However, there is a subtlety here: our affine scheme comes equipped with a bifiltration, and while $X$ is isomorphic to $X_{V}$ as schemes, they may not be isomorphic as bifiltered schemes. One important open question is to understand the corresponding bifiltration structure on $X_{V}$. This is especially prominent for theories with non-trivial Higgs branch, which we discuss in Section 5, where we find it important to add to the ring $R$ an additional generator, beyond the strong generators of the VOA, and interpreted as corresponding to the stress-tensor.
@@ 2.1 Arc spaces: a primer
The jet scheme $J_{n}(X)$ admits a natural bigrading defined by
@@
Practically, this involves enumerating the infinitely many coordinates and polynomials in the ideal for the arc space of a given scheme $X$. Intuitively, the arc space captures formal smooth paths (or arcs) based at points of $X$, with the bidegree keeping track of derivative order. For example, the first jet scheme $J_{1}(X)$ can be identified with the total space of the tangent bundle of $X$.
@@ 2.2 Vanishing short multiplets in Argyres–Douglas theories
This phenomenon can be understood from the perspective of the associated VOA. The VOA corresponding to the $(A_{1},A_{2n})$ theory is the Virasoro minimal model $M(2,2n+3)$, which contains a null state of the form $((L_{-2})^{n+1}+\cdots)\Omega$ at level $2n$, where $\Omega$ denotes the vacuum state. This reflects the vanishing of the $\widehat{\mathcal{C}}_{n(\frac{n}{2},\frac{n}{2})}$ multiplet in the physical theory.
@@
where $\operatorname{PE}$ denotes the standard plethystic exponential. The term $\frac{q^{2}}{1-q}$ arises from the contribution of the stress tensor, while the term $\frac{q^{2n+2}}{1-q}$ reflects the absence of the $\widehat{\mathcal{C}}_{n(\frac{n}{2},\frac{n}{2})}$ multiplet. [^10] The Schur index alone does not uniquely determine which short multiplet is absent; this requires additional input from OPE selection rules [67, 71, 63, 56]. In some cases, further information can be obtained by studying the Macdonald or full superconformal index, combined with character decomposition of short multiplets [2, 78].
@@
For low values of $N$, this motivates us to define the affine schemes as follows:
$(A_{2},A_{3}):$ $\quad x^{2}+y^{3}=0\,,\ xy=0\,$, (2.18a)
$(A_{2},A_{4}):$ $\quad x^{2}+y^{3}=0\,,\ xy^{2}=0\,.$ (2.18b)
Here $x$ and $y$ comes from $W$ and $T$ respectively. Importantly, as there are few monomials that can contribute at the relevant degrees, any coefficients can be absorbed by linear redefinitions of the coordinates. Thus, for $N=4,5$, we do not need to concern ourselves with the coefficients; for $N>5$, this is no longer the case. For $N=7,8$, the coefficients appearing in the polynomials at each degree can be fixed (up to linear redefinitions) by demanding that the Hilbert series of the arc space reproduces the Schur index. We find
$(A_{2},A_{6}):$ $\quad x^{3}+4xy^{3}=0\,,\ 3x^{2}y+y^{4}=0\,$, (2.19a)
$(A_{2},A_{7}):$ $\quad y^{5}+5y^{2}x^{2}=0\,,\ 2x^{3}+3xy^{3}=0\,.$ (2.19b)
In each of these cases, the scheme reduces to a point when reduced to an affine variety. But, as schemes, they have a more intricate singularity structure, which captures the physics of the indices.
We can further generalize this to an arbitrary $(A_{k-1},A_{N-1})$ theory. The associated VOA is a $W_{k}$-algebra, containing k1k-1 generators of spins $2,3,\cdots,k$. The Schur index (or the vacuum character of the $W_{k}$-algebra) is given by
@@
[…]
where we ignored the coefficients. Writing $(W_{2},W_{3},W_{4})\to(z,y,x)$ will give the desired relations for the point-like singularity in three dimensions. It is possible to get the same relations from $F=z^{5}+z^{3}x+zx^{2}+z^{2}y^{2}+y^{2}x$ and taking derivatives with respect to x,y,zx,y,z. [^12] More generally, it is conjectured [84] that for the Argyres–Douglas theories of type $(A_{k-1},A_{n-1})$ with $(k,n)=1$, Zhu’s $C_{2}$-algebra $R_{V}$ is given by a simple Jacobi algebra of the form $R_{V}=\mathbb{C}[w_{1},\ldots,w_{\ell}]\Big/\left(\frac{\partial F}{\partial w_{1}},\ldots\frac{\partial F}{\partial w_{\ell}}\right)\,$, where $F=F(w_{1},\ldots w_{\ell})$ is the sum of monomials of fixed degree and the degrees for $w_{i}$ are given as $i+1$. However, the numerical coefficients appearing in the polynomials that define the affine scheme affect the Hilbert series of the arc space, and therefore it is necessary to be precise about the coefficients before attempting to match the indices. In general, the Jacobi algebra which is conjectured to be identical with Zhu’s $C_{2}$-algebra does not produce the correct Schur index.